​Madeoff's Mortgage. Bernie Madeoff pays ​$220 comma 000 for a new​ four-bedroom 2,400-square-foot home outside​ Tonopah, Nevada. He plans to make a 20​% down​ payment, but is having trouble deciding whether he wants a 15​-year fixed rate ​(6.398​%) or a 30​-year fixed rate ​(6.873​%) mortgage. a. What is the monthly payment for both the 15​- and 30​-year ​mortgages, assuming a fully amortizing loan of equal payments for the life of the​ mortgage? b. Assume that instead of making a 20​% down​ payment, he makes a 10​% down​ payment, and finances the remainder at 7.124​% fixed interest for 15 years. What is his monthly​ payment? c. Assume that the​ home's total value falls by​ 25%. If Bernie sells the house at the new market​ value, what would be his gain or loss on the home and​ mortgage, assuming all of the mortgage principal​ remains? Use the same assumptions as in part a. a. What is the monthly payment for both the 15​- and 30​-year ​mortgages, assuming a fully amortizing loan of equal payments for the life of the​ mortgage? The monthly payment for the​ 15-year mortgage is ​$ nothing.
Answer #1
A) Home Value Down Payment (20%) Total Loan (80%)
$220,000.00 $44,000.00 $176,000.00
For a 15 Year Loan, total instalments would be = 15 x 12 = 180
For a 30 Year Loan, total instalments would be = 30 x 12 = 360
Rate for 15 Year loan is 6.398% per annum. Converting it to monthly rate = 6.398/12 = 0.53%
Rate for 30 Year loan is 6.873% per annum. Converting it to monthly rate = 6.873/12 = 0.57%

The Formula Used is:

12 A = Pez ( 1. A - Amortization P - Principal 1 - Interest Rate n - Term (expressed in months)

Where P = $176000
i=0.0053, n=180 (for 15 year loan)
i=0.0057, n=360 (for 30 year loan)
Plugging in these values we will get
Total Monthly Payment for a 15 Year Fully Amortizing Mortgage loan is $1,523.22
Total Monthly Payment for a 30 Year Fully Amortizing Mortgage loan is $1,155.95
B) Home Value Down Payment (10%) Total Loan (90%)
$220,000.00 $22,000.00 $198,000.00
For a 15 Year Loan, total instalments would be = 15 x 12 = 180
Rate for 15 Year loan is 7.124% per annum. Converting it to monthly rate = 7.124/12 = 0.59.%

The Formula Used is:

Where P = $198000
i=0.0059, n=180
Plugging in these values we will get
Total Monthly Payment for a 15 Year Fully Amortizing Mortgage loan is $1,793.47